### v1

# Geometric Reasoning in the Embedding Space   Jan Hůla1,2, David Mojžíšek2, Jiří Janeček2, David Herel1, Mikoláš Janota1

Abstract
In this contribution, we demonstrate that Graph Neural Networks and Transformers can learn to reason about geometric constraints. We train them to predict spatial position of points in a discrete 2D grid from a set of constraints that uniquely describe hidden figures containing these points. Both models are able to predict the position of points and interestingly, they form the hidden figures described by the input constraints in the embedding space during the reasoning process. Our analysis shows that both models recover the grid structure during training so that the embeddings corresponding to the points within the grid organize themselves in a 2D subspace and reflect the neighborhood structure of the grid. We also show that the Graph Neural Network we design for the task performs significantly better than the Transformer and is also easier to scale.
Keywords: Machine Learning, Geometric Reasoning, Neural Networks, Embedding Space
1 Introduction
How do Neural Networks (NNs) solve problems that require spatial reasoning? Many papers already demonstrated that autoregressive Transformers, e.g. Momennejad et al. (2023); Wu et al. (2024); Yamada et al. (2023), and Graph Neural Networks (GNNs), e.g. Li et al. (2023); Teodorescu et al. (2022), are able to learn to solve such problems but it is not clear what is the mechanism they discovered.
A notable example is the work by Trinh et al. (2024) who developed a system named AlphaGeometry, which was able to solve problems that appeared on the International Mathematical Olympiad. The system contains an autoregressive Transformer that takes as input a sequence of tokens describing the problem in the language of geometric relations and is trained to predict auxiliary points useful for finding a proof for the given statement. Unfortunately, there is no explanation of the process by which the model manages to propose useful auxiliary points.
A human dealing with such a problem would try to form a mental image of the construction used in the problem (most likely by first drawing it). We can ask a natural question whether NNs could also form such “mental image”, which would reflect the spatial configuration of points described in the problem. We can also ask whether an autoregressive Transformer is a suitable model for such a problem and whether a GNN, which eliminates a lot of symmetries [^1] Variable renaming and constraint reordering., could be easier to train and be more scalable.
In this contribution, we take a closer look at these questions in the domain of geometric Constraint Satisfaction Problems (CSPs). In order to simplify the problem as much as possible, we create a simple CSP language with several geometric constraints (relations) for which the domain is a set of points in a discrete 2D grid of certain size. Each instance of this CSP uniquely describes a hidden figure whose points are the solution of the instance. As the discrete grid is finite, we can assign a token/class [^2] When using a Transformer, each point is represented by a token; when using GNN, each point corresponds to a class. to each point and train the model to predict the points in the hidden figure with a cross-entropy loss.
Our analysis reveals interesting properties of the learned model. After visualizing a low-dimensional projection of the embeddings corresponding to individual points, we can see that they organized themselves in a 2D grid which they represent.
We also show that during inference, the embeddings of unknown variables evolve into a configuration that reflects the hidden figure described by the problem. Finally, we show that a GNN designed for the CSP performs significantly better than the autoregressive Transformer and is also able to scale to larger grids. Even though the prediction task we deal with is much simpler than the one required in systems such as AlphaGeometry, we believe that our findings provide useful insights for designing and thinking about models used for geometric reasoning.
The rest of the paper has the following structure. In Section 2, we introduce related work. Section 3 contains the description of our experimental setup. This includes the two types of architectures and the process for the generation of problems. In Section 4, we describe results of experiments for both models and provide an analysis of failure modes and scaling laws (in 4.3 and 4.4, respectively) for different sizes of the 2D grid. We test how well the described methods scale with changes in data. We discuss the results and limitations of our work in Section 5. Supplementary materials can be found in Appendix.
2 Related Work
Reasoning About Geometry
Geometric reasoning has been a focus of research for decades Wen-Tsun (1986), with Wu’s method Wu (2008) considered state-of-the-art until recently. In 2024, AlphaGeometry Trinh et al. (2024) surpassed Wu’s method on International Mathematical Olympiad problems, utilizing a combination of symbolic deduction and neural language model which predicts auxiliary constructions based on problem statement which is described using geometric relations. Our work is partially inspired by AlphaGeometry but we study much easier setup in which the model predicts points in a 2D grid that satisfy the relations used to describe the construction.
We were also partially inspired by the work of Hůla et al. (2024) which provides explanation for the process by which GNNs can learn to solve Boolean CSPs, concretely deciding satisfiability of Boolean formulas. We augment their setup for the domain of geometric CSP. In their work, they show an empirical evidence that the GNN learns to act as a first-order solver of a semidefinite programming (SDP) relaxation of MAX-SAT. MAX-SAT is an optimization version of SAT and in the SDP relaxation, vectors assigned to variables are being optimized in order to minimize an objective which captures how many constraints are satisfied.
In the domain of geometric CSPs, Krueger et al. (2021) developed an algorithm for automatic construction of diagrams for geometry problems. The algorithm works by optimizing coordinates of points by gradient descent to maximize an objective function which captures how well are the geometric constraints of the problem satisfied. We emphasize that their algorithm was manually constructed and not learned.
Spatial Reasoning in Neural Networks
Another related area of research is focused on the ability of NNs to do spatial reasoning in a broader sense, e.g. spatial navigation, planning and interaction with physical objects.
Several papers explore spatial reasoning in the context of Large Language Models (LLMs) such as ChatGPT or LLama Touvron et al. (2023). In Momennejad et al. (2023) and Yamada et al. (2023), they point out various failure modes in a planning and navigation capability of large pool of LLMs. Wu et al. (2024) developed a prompting technique called Visualization-of-Thought (VoT), which significantly improved LLM performance in visual navigation tasks. Most similar to our work is the work by Ivanitskiy et al. (2023) who probe small Transformer models trained for maze navigation in order to understand the representation it uses to solve the task. Unlike the previously mentioned works, Janner et al. (2018) uses reinforcement learning to train a language model based on a recurrent NN (RNN) for text-based spatial reasoning.
Apart from language models, other papers also explore spatial reasoning in the context of GNNs. Li et al. (2023) developed a GNN for spatial navigation, which utilizes a memory mechanism for long-range information storage and Teodorescu et al. (2022) introduced the SpatialSim benchmark together with a GNN that is trained to recognize spatial configurations. Our work contributes into this area by partially elucidating the mechanism by which NNs are able to reason about geometric relations and by showing that GNNs which use the graph of relations/constraints and variables are much easier to train and scale.
3 Experimental Setup
To test the ability to reason about geometric relations, we develop a simple generative procedure of problems with solutions that allows to control the difficulty and size of the problem. We test two types of models, GNNs and Transformers, on data generated with various generator settings.
3.1 Data Generation
Our goal was to create the simplest possible synthetic dataset that would enable us to study the mechanism which NNs use to reason about geometric relations. We wanted to know whether the geometry of embeddings within the network somehow reflects the geometric properties of the problem. In the task of auxiliary point prediction for which AlphaGeometry was using a language model, it is, for example, not clear whether the model is taking the geometric nature of the constraints into account or treats the language of the problem abstractly without any metric interpretation/grounding.
Therefore, we designed a generator for CSPs whose solutions can be interpreted as hidden figures described by the input constraints. The task of the model is to predict positions of the points mentioned in the constraints. We simplify the problem so that the hidden figures lie on a discrete grid, and we can therefore treat the prediction of the positions as a classification task. We note that one can also treat the task as a regression and train the model to process figures with arbitrary 2D coordinates.
Our CSP is using four types of constraints: M, R, S and T. The constraint $M(a,b,c)$ says that the point $b$ is a midpoint of the line segment given by $a$ and $c$ and $R(a,b,c,d)$ says that points $a$, $b$ give an axis of symmetry around which $c$ and $d$ reflect each other. The constraint $S(a,b,c,d)$ says that points $a$, $b$, $c$, $d$ form a square and $T(a,b,c,d)$ says that the vector $d-c$ is a translation of the vector $b-a$. In order to have unique solutions, we also need to fix several points to concrete positions. We can achieve that by additional constraint P for which $P(a,[1,1])$ signifies that point $a$ lies at position $[1,1]$. For each constraint, certain number of variables are required to be known so that the constraint can be uniquely resolved, here we call them determining variables. Once these variables are fixed, the rest of variables is uniquely determined. These are called dependent variables. For the constraint $M$, $2$ variables are determining and $1$ is dependent, i.e. if we know positions of $2$ variables, the last one could be resolved. For the constraint $R$, $3$ variables are determining and $1$ is dependent. For the constraint $S$, $2$ variables are determining and $2$ are dependent (i.e., we assume a fixed spatial ordering of the square vertices). For the constraint $T$, $3$ variables are determining and $1$ is dependent. The last subplot in Figure 2 shows a visualization of the following instance of the CSP:
$T(F,D,E,G)\wedge T(C,G,A,H)\ \wedge$
$S(H,I,G,J)\wedge S(B,J,L,K)\ \wedge$
$P(A,[6,15])\wedge P(B,[12,16])\wedge P(C,[12,14])\ \wedge$
$P(D,[14,12])\wedge P(E,[12,14])\wedge P(F,[12,9])\ ,$
where $A-F$ are known points and $G-L$ are unknown.
We generate the problem instances in such a way that the correct assignment to all variables can be obtained by incrementally resolving constraints that can be uniquely resolved. Therefore, to sample a problem, we sample a random directed acyclic graph (DAG) with nodes corresponding to types of constraints and edges corresponding to dependencies. Then we add variables to the constraints so that each constraint within graph can be uniquely resolved given that the parent constrains are already resolved (the parent constraints therefore need to contain the determining variables). To produce a problem with a unique solution, we fix the required number of variables appearing in each of the root constraints which have no parents (i.e. arbitrary 2 variables in the S constraint, 3 variables for the T constraint, etc.) We use the Z3 solver De Moura and Bjørner (2008) to obtain the assignment for all other points. If the resulting figure is larger than the size of the grid, we reject the problem, otherwise we place it to a random position within the grid. The constraints with the fixed points are given as an input to the model, the positions of the remaining variables are given as labels.
3.2 Models
In our experiments, we tested two types of architectures: GNNs and autoregressive Transformers.
Graph Neural Network
The architecture of the GNN is inspired by the model presented in Hůla et al. (2024). It is a GNN operating on a bipartite graph with $n$ variables and $m$ constraints defined by Equations (1) and (2) below which are applied recurrently for several rounds. Equation (1) updates the embeddings of variables which are, after the last update, used to classify a given variable to a point within the grid. This is achieved by applying a linear layer $L(\cdot)$ on each variable embedding. As mentioned in 3.1, some variables need to be fixed to concrete points so that the problem has a unique solution. The embeddings of these variables are initialized with the embedding layer whose weights are shared with the final classification layer $L(\cdot)$ [^3] Similarly as token embeddings share weights with the classification head in language models.. They are also not updated during message passing. The embeddings of variables that are not fixed are initialized by random unit vectors and updated after each message passing step. The update equations have the following form:
$X^{t+1}$ $=U_{X}\left(A_{X}Z^{t},X^{t},\Phi_{X}\right)\,,$ (1)
$Z^{t}_{c}$ $=U_{c}\left(A_{c}X^{t},Z^{t-1},\Phi_{c}\right)$ $\forall c\in C\,,$ (2)
where $X^{t}\in\mathbb{R}^{n\times d}$, $Z^{t}\in\mathbb{R}^{m\times d}$, $A_{X}\in\mathbb{R}^{n\times m}$, $A_{c}\in\mathbb{R}^{4m\times n}$ are matrices and $C=\{M,R,S,T\}$ is the set of constraint types. Equation (1) updates the embeddings of variables and Equation (2) updates the embeddings of the constraints. The latter equation is further indexed by $c$ which reflects the fact that we are using multiple types of constraints which require distinct update functions. The update functions $U_{x}$ and $U_{c}$ are in our case realized as LSTMs (Hochreiter, 1997) parameterized by $\Phi_{X}$ and $\Phi_{c}$, respectively. Technically, each LSTM also updates a cell state which is not mentioned in the equations. These LSTMs can be also replaced by simple RNNs but we found the LSTMs easier to optimize.
The update of each embedding happens independently of the other embeddings, i.e. in Equation (1), each embedding of a variable (row in the matrix $X^{t}$) is updated independently by the same LSTM which takes as input the aggregated message from the constraints containing this particular variable. The aggregated message is simply the sum of the relevant constraint embeddings and it is realized by multiplying the constraint embedding matrix $Z^{t}$ by the matrix $A_{X}$. The LSTM $U_{c}$ which updates the constraint embeddings, differs by the fact that the aggregated message is not obtained as a sum but as a concatenation of variable embeddings. The embeddings are concatenated in the order in which the variables appeared within the constraints. For example, for the constraint $S(a,b,c,d)$, we concatenate the embeddings of variables $a,b,c,d$ in that order. One can intuitively view the embeddings of variables as if they are representing the values of these variables (points) and embeddings of constraints as if they represent the information of what is needed to satisfy the constraint. For this reason, the function which updates the constraint embeddings cannot be permutation invariant because different order of the determining variables results in different values for the dependent variables. The embdding dimension $d$ is set to $96$ for the experiment in 4.2 and $256$ for the experiment in 4.4. The number of iterations of the model is set to $20$. Other hyperparameters for training the GNN can be found in Table 3.
Autoregressive Transformer
The autoregressive Transformer model is based on the GPT-2 architecture developed by Radford et al. (2019) with rotary embeddings Su et al. (2024). It takes as input a sequence of tokens representing the problem together with a query for a variable we want to predict. For example, the input with just one constraint could have the following form: S ( [0,0] [0,1] C D ) ? D. It is querying the position for variable D within a square with two known points. The model reads this sequence of tokens (where the positions of points such as [0,0] correspond to one token) and is trained to predict the token which corresponds to the the correct position of variable D ([1,0] in this case). This means that from each CSP produced by the generator described in Section 3.1, we extract one sequence for each unknown point. After experimenting with the hyperparameters of the model, we set the number of layers to 6, number of heads to 6 and the embedding dimension to 256. We also experimented with a recurrent application of one layer, as done in Dehghani et al. (2018), to more closely mimic the GNN, but having separate weights for each layer produced better results. Other hyperparameters of the model and for training can be found in Table 3.
Table 1: Performance of the GNN when allowing different types of geometric constraints in problems. We report both point-wise accuracy (proportion of correctly predicted points) and complete accuracy (proportion of problems where all points were correctly predicted). Each row shows results for different combination of constraint types: SQUARE (S), TRANSLATION (T), MIDPOINT (M) and REFLECTION (R). The distribution of sampling a each type of constraint is uniform. All problems are for a grid size $20\times 20$.
Point Accuracy Complete Accuracy
S and T 98.4% 84.7%
S, T and M 89.5% 71.8%
S, T, M and R 67.7% 38%
4 Experiments
4.1 Comparison of the Two Architectures
To compare the two architectures, we measure the accuracy of individual point prediction (point accuracy). In later experiments, we also report the accuracy in terms of correctly assigning all variables within the problem (complete accuracy).
The experiments on the synthetically generated data show that the GNN performs significantly better than the Transformer, as one could expect. In 4.4, we show that we were able to train the GNN on grid sizes up to $80\times 80$ points to a validation accuracy larger than $90\,\%$. In comparison, the Transformer achieved accuracy of $90\,\%$ only if we train it on a grid size $10\times 10$ and limit the number of constraints to $2$. If we train the model on a more complex setting with the grid size $20\times 20$ and up to $6$ constraints, it reaches an accuracy of approximately $30\,\%$.
For completeness, we also trained the Transformer model to find the assignments to variables using Chain-of-Thought (CoT) Wei et al. (2022) by imitating a log from a simple solver. The solver resolves the constraints one by one in the topological order of the DAG of constraints mentioned in Section 3.1. Using this way of training, the Transformer learns to predict the positions of variables within $20\times 20$ with approximately 50% point accuracy. We did not explore this direction further as reasoning with a CoT is orthogonal to reasoning in the embedding space on which we focus in this work. More details about the CoT experiment can be found in Appendix F.
Given the superior results achieved by the GNN, we conduct the main analysis of the embeddings on the GNN. Appendix E contains similar analysis and visualization for the Transformer (see Figure 7).
We also note that the model accuracy can be further improved by leveraging multiple prediction attempts since incorrect predictions often fall close to their true positions (as shown in Appendix D) and according to preliminary tests, sampling multiple different initial embeddings increases the chance of predicting the correct assignment.
4.2 Visualizing the Embeddings of Individual Points
In the following text, we use the terms static embeddings and dynamic embeddings. By static embeddings we mean the embeddings of points of the grid which are used for initialization of known points and are shared with the classification head of both models. By dynamic embeddings we mean the embeddings of the unknown variables which are updated throughout the forward process.
Both types of models are trained to predict positions of unknown points within the instance. This is done by a linear layer which computes the logits for each point. As already mentioned, the weights of this layer are shared with the embedding layer representing the position of known points.
When visualizing low-dimensional projection of the static embeddings corresponding to individual points, we found that they organize themselves into a 2D grid they represent. In Figure 1, we show how this organization emerges during training and how it is connected to the precision of the prediction (three dimensional projection is depicted in Appendix C).
Figure 1: Evolution of static embeddings throughout training. Lines and colors represent the connectivity of points within the 30$\times$30 grid in one direction. UMAP McInnes et al. (2018) is used as a projection method from the original 96 dimensions of the embedding space. We found that UMAP works best for larger grids; however, simpler PCA projections worked better for smaller grids (15$\times$15 and less points).
We stress that the existence of grid structure of the data domain and the existence of geometric figures related to constraints is given to a model only indirectly through the constraints and their solutions. The whole training dataset can be thought of as a set of constraints written in an unknown language, and the goal of training is to find a model for this language which will enable correct prediction. As the prediction is based on the similarity (given by inner product) of the dynamic embeddings of variables and static embeddings of points, it is not that unexpected that the discovered model reflects the geometry behind this language.
4.2.1 Solution Process
In order to find the solution, the GNN model iteratively moves dynamic embeddings in a high-dimensional space. During the forward pass, the same update rule is applied over and over again. Hence, it is possible to extract the embeddings in each step of the model to get a better understanding of the underlying process.
Figure 2 shows an example of the solution process for a random problem given to the GNN. After each update, the closest static point embedding is taken for each variable of a problem and is visualized as the corresponding grid position.
Note that in this example, the GNN first finds a close approximation to the hidden configuration and then refines it. Here, the squares are first “approximated” by quadrilaterals which then converge to exact squares. It can also be observed that one square constraint is approximately satisfied earlier than the other which reflects the fact that the other constraint can be resolved only after the first constraint is resolved (as explained in the next section).
Figure 2: Visualization of the solution process. The red points A, B, C, D, E, F (C = E) are known, and the blue points G, H, I, J, K, L need to be predicted. There are two translation constraints: T (F, D, C=E, G), T (C=E, G, A, H) and two square constraints: S (H, I, G, J), S (B, J, L, K). Translations are marked by a dotted line and squares by a solid line. The network is trained to predict the result in $15$ iterations, of which initial state and results after iterations $3,5,7,9,11,12,13$ were chosen for illustration. The network gradually improves the result over the iterations: the first translation with only one unknown point G is solved, followed by finding the point H of the second translation. After translations, both squares are solved.
4.3 Analysis of Incorrectly Classified Points
(a) Relative incorrect point distribution.
(b) Total point distribution.
Figure 3: Predicted points classified based on the number of resolving steps (on the constraint level) that are
necessary to determine the point positions.
To understand the failure modes of the GNN, we analyze predictions of approximatelly 10k instances of our CSP. These instances contain around 29k points to be classified. As explained in Section 3.1, the constraints of each instance could be organized into a DAG by a dependency given by the variables. A constraint A depends on a constraint B if the determining variables of A are the dependent variables of B. To resolve a given variable, we can start with the constraint which contains it and recursively resolve all the constraints on which this constraint depends. Once all these constraints are resolved, we can also resolve the constraint in which this variable appears, and thus we get its value. The constraints which need to be resolved in order to resolve a given variable will constitute a sugraph of the DAG and the number of constraints in this subgraph determines the number of steps which are needed to resolve the given variable. If the variable is contained in multiple constraints, then we can consider the one which requires the least number of resolving steps.
We hypothesized that the incorrect predictions will be correlated with the number of resolving steps required for a given variable. The barplot in Figure 3a confirms this hypothesis. It shows that the relative failure rate (incorrect predictions vs. all predictions for a particular point) depends almost monotonically on the number of resolving steps required. Monotonicity of this dependency is violated only in the case of 7 resolving steps which could be attributed to the insufficient number of tested points to robustly estimate the failure rate. The absolute numbers of predicted points in the generated problems are shown in Figure 3b. We also mention that the incorrectly classified points are very often classified to a point lying very close the correct point (see Appendix D).
4.4 Scaling the Size of the Grid
To get a sense of how the sample complexity depends on the size of the grid, we train several models on different sizes of the grid and different amount of training samples. To achieve faster training, we conducted these experiments with problems which contained only two types of constraints (S and T).
The problem generator mentioned in 3.1 produces problems which have on average around four constraints. Both types of constraints ($S$ and $T$) are sampled with equal probability. Therefore, we can assume that an average problem has two constraints of type $S$ (which is determined by $2$ points) and two constraints of type $T$ (which is determined by $3$ points). If we denote the number of points on the side of the grid by $n$, then we can estimate the number of unique problems in the grid of size $n\times n$ to be $n^{20}$. There are $n^{2}$ possible point positions and we independently sample 6 points for two constraints of type T and 4 points for two constraints of type S, together yielding $(n^{2})^{10}$ possibilities each of which determines one instance. This number should be viewed as an upper bound because it ignores the fact the constraints can share variables.
To test the dependence of the validation accuracy on the grid size and number of training samples, we use the following values of $n$: $10$, $20$, $30$, $40$, $50$, $60$, $70$, $80$. This results in the following number [^4] After rounding the exponent to the nearest integer. of possible problems for each $n$, respectively: $10^{20}$, $10^{26}$, $10^{29}$, $10^{32}$, $10^{34}$, $10^{35}$, $10^{37}$, $10^{38}$.
The sizes of the training set for each grid size are in the range from $5$k to $800$k. The relationship between the validation accuracy, the grid size and the size of the training set can be seen in Figure 4a. In Figure 4b, we plot the relationship between the grid size and sizes of the training set for which the validation accuracy exceeded $90\,\%$. As can be seen, in this range of grid sizes, the sample complexity grows much more slowly than the number of possible examples.
4.5 Impact of Constraint Types
We also analyzed how model performance changes when incorporating additional geometric constraint types while maintaining consistent training conditions on a $20\times 20$ grid and a fixed training set size of 70k examples (Table 1). Adding each new constraint type (M, R) to the base set (S, T) led to substantial decreases in both point-wise and complete-figure accuracy. The observed decrease in accuracy appears steeper than would be expected from linear scaling with constraint count, suggesting increased problem complexity. Given our findings on scaling laws in Section 4.4, it is likely that higher accuracy could be achieved with larger training datasets.
(a) Accuracy vs number of training samples for
different grid sizes.
(b) Grid size vs sample size required for accuracy
$>90\,\%$.
Figure 4: Scaling laws for different sizes of the 2D grid. For a grid of size $n$, there are $n^{2}$ points.
5 Limitations and Discussion
We note that we study a much simpler setup than was studied in AlphaGeometry. Predicting positions of points satisfying a set of geometric constraints is straightforward in comparison with predicting useful auxiliary points. Also, in our case, the unknown points need to lie on a discrete 2D grid so that we can train the model with a classification loss. The simplified setup turned out to be sufficient to study how NNs can reason about geometric constraints. Training the model for predicting auxiliary points requires a large computational budget and, moreover, the authors did not release the training dataset, which is also expensive to create. We also mention that we use only four types of constraints, but the whole setup could be easily extended to different sets of constraints by adding more update functions in Equation (2). These simplifications enabled us to provide partial understanding of the process by which the tested models are able to solve individual problems and to demonstrate the benefits of using a GNN instead of a Transformer.
Compared to the GNN which Hůla et al. (2024) used for Boolean satisfiability, our GNN can deal with fewer constraints and points. In their case, the GNN was able to handle hundreds of constraints, whereas our GNN struggles with instances that contain more than 10 constraints. This could be caused by the fact that the domain of our constraints is much larger, but it could also be the case that different architecture of the GNN (with different update functions) could scale to larger problems. We leave this exploration for future work.
It is also possible that a different training procedure might result in significantly easier training. The inference process depicted in Figure 2 and the findings of Hůla et al. (2024) suggest that the GNN could implicitly optimize an unknown objective during the iterative forward pass. Finding an expression for this objective is another possible direction for future work. It could be used to train the network in an unsupervised way which should be much easier that training it with the classification loss. Another obvious direction is to train the model as a diffusion model which “denoises” randomly assigned points to points of the solution.
Lastly, we mention that our experimental setup can be extended to more complex CSPs which could contain temporal relations and relations between various entities. We believe that such CSPs could be very useful for studying how NNs can generalize to unseen situations Abbe et al. (2023) and how they can discover models of the world without grounding Wong et al. (2023). Studying geometric CSPs has the advantage that the domain has an obvious geometric interpretation which is visible in the embedding space.
6 Conclusion
We have shown that GNNs as well as Transofmers can learn to solve geometric CSPs and provided several insights into the process by which they find the solution. The visualizations show that when processing the problem, models form the hidden spatial configurations described by the problem in the embedding space. During training, the static embeddings of individual points in the 2D grid organize themselves within a 2D subspace and reflect the neighborhood structure of the grid. We showed that the occurence of errors depends on the number of steps required to resolve a given variable and the number of types of constraints present within the CSP language. Lastly, we showed that GNNs are much easier to train and can be scaled to significantly larger grids than Transformers.
Appendix A Appendix
Appendix B Model and Training Hyperparameters
The final hyperparameters for both model architectures and for their training can be seen in Tables 3 and 3.
Table 2: Model hyperparameters and training settings for the GNN.
Graph Neural Network
Hyper-parameter Value
Embedding dimension 96,256
Embedding update iterations 20
Batch size 256
Number of epochs 200
Optimizer AdamW
Initial learning rate $10^{-3}$
Warmup steps 5
Learning rate scheduler CosineAnnealingLR
Gradient clipping norm $0.6$
Weight Decay $10^{-6}$
Table 3: Model hyperparameters and training settings for the Transformer.
Transformer
Hyper-parameter Value
Number of layers 6
Number of heads 6
Embedding dimension 256
Positional embeddings (RoPE)
Batch size 512
Number of epochs 200
Optimizer AdamW
Learning rate $5\times 10^{-4}$
Warmup steps 200
Learning rate scheduler Linear
Gradient accumulation steps 1
Gradient clipping norm 1.0
Evaluation steps 10
Special tokens [SEP], [UNK],
[PAD], [MASK]
Appendix C Embedding Visualization for the GNN in 3D
As mentioned in 4.2, the static embeddings of individual points organized themselves into a grid which they represent. In Figure 5, we visualize the 3D projection of these embeddings, over the course of training. At the beginning, the randomly initialized embeddings form a single spherical cluster, which later unfolds into a U-shape surface and finally into a flat 2D surface.
Figure 5: We used UMAP to show that the internal structure of static embeddings discovered by the model reflects the 2D essence of the grid. The visualization shows a projection into 3D taken after 1, 5, 15, 50 and 100 epochs of training.
Appendix D Distribution of the Error Magnitudes
The histogram in Figure 6 depicts a distribution of Euclidean distances between the incorrectly predicted point position and the ground truth position on the grid. As can be seen, most incorrectly predicted positions lie very close to the ground truth position.
Figure 6: Histogram of Euclidean distances of incorrectly predicted points by the GNN, from the ground-truth points. Distances with low number of occurrences ($<5$) are grouped together (bar “Others”). As can be seen, most incorectly predicted points were close to the ground-truth points.
Appendix E Embedding Visualization for the Transformer Model
In Figure 7, we show the grid-like structure of static point embeddings discovered by the Transformer. In comparison with Figure 5 which shows the emergence of the corresponding structure for GNN, the grid is now of smaller size, namely 10$\times$10, since the Transformer was not able to scale into larger sizes with sufficient accuracy as discussed in 4.1.
Figure 7: Projection into 2D using UMAP illustrates that static token embeddings of the Transformer self-organized into a grid-like structure which the visualized tokens represent. The edges show the connectivity of the points in the original grid.
Appendix F Chain-of-Thought Training
In order to train the Transformer to produce a chain-of-thought which assigns the variables incrementally, we implemented a simple solver which logs its steps and the resulting logs are used for imitation. The solver first orders the constraints according to a DAG mentioned in Section 3.1 and then resolves the constraints one by one, starting from the root constraints. As the solver traverses the DAG, it logs the constraint of a given node, the values of already assigned variables within the constraint and lastly the computed values for the remaining variables. We also include few keywords into the log which delimit the provided information. Example of the log for a random problem is shown below:
TRANSLATION ( 1 0 2 3 ) , TRANSLATION ( 5 4 7 6 ) , SQUARE ( 8 7 9 3 ) , SQUARE ( 11 8 10 3 ) , TRANSLATION ( 8 7 12 3 ) ; fixed 0 = #696 , 1 = #617 , 2 = #978 , 4 = #577 , 5 = #498 , 6 = #731 ; Solution begins ; Con TRANSLATION ( 5 4 7 6 ) ; Known 5 = #498 , 4 = #577 , 6 = #731 ; Impl 7 = #652 ; Con TRANSLATION ( 1 0 2 3 ) ; Known 1 = #617 , 0 = #696 , 2 = #978 ; Impl 3 = #1057 ; Con SQUARE ( 8 7 9 3 ) ; Known 7 = #652 , 3 = #1057 ; Impl 8 = #462 , 9 = #867 ; Con SQUARE ( 11 8 10 3 ) ; Known 8 = #462 , 3 = #1057 ; Impl 11 = #247 , 10 = #842 ; Con TRANSLATION ( 8 7 12 3 ) ; Known 8 = #462 , 7 = #652 , 3 = #1057 ; Impl 12 = #867 ; Solution ends
Variables are expressed by a number (1, 2, 3, etc.) and individual points are expressed with point ID with a # symbol in front (#696, #617, etc.). The input has two parts, a problem statement and a solution. We train on the whole input, but exclude the problem statement from the computation of the loss. For validation, we include only the problem statement. The keyword Con marks the selected constraint, Known marks the known variables which appear in the selected constraint with their values and Impl marks the newly deduced variables with their values.